OpenStudy (anonymous):

First factor the trinomial as a binomial square and then use the square root method to solve. Round the solution to the nearest tenth. x^2 - 14x + 49 = 3 A. {5.3, 8.7} B. {4.8, 9.2} C. {3.5, 7.5} D. {6.2, 8.3}

OpenStudy (anonymous):

Factoring: binomial squares have certain rules they follow: for instance, (a+b)^2 = a^2 + 2ab +b^2 and (a-b)^2 = a^2 - 2ab + b^2. This problem looks like it follows the second rule. See if you can find a and b. Forget about the 3 when you do this

OpenStudy (anonymous):

Ok creative you lost me

OpenStudy (anonymous):

haha ok, my bad. so basically, if you let x=a, you can relate my equation to (x-b)^2 = x^2 -14x +49. All you need to do now is figure out what b is. we know that the last term (49) is b^2. So you can just solve b^2=49 to find b

OpenStudy (anonymous):

You follow? Square Root Method: once you get your b, your new factored equation will look something like this: (x-b)^2 = 3 (only, with a number in b) The square root method just means you take the square root of both sides so you get \[(x-b) = \sqrt{3}\]Then solve for x (because you should have gotten your b earlier)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

did you get b?

OpenStudy (anonymous):

-14x

OpenStudy (anonymous):

um, no not the term....i should have used different variables, sorry. From my equation b^2=49 b is 7 (plus or minus).

OpenStudy (anonymous):

:/ lost me again/

OpenStudy (anonymous):

ok, that's probably my fault. sorry!! do you know what i was saying in my 2nd post?

OpenStudy (anonymous):

No not really

OpenStudy (anonymous):

Did you get the bit about the (a-b)^2 rule?

OpenStudy (anonymous):

I would tell you just to use the quadratic formula to factor the equation, but the question doesn't want you to... it would have been easier to explain.

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Ok, what i did was to try to transform the given equation into that (a-b)^2 rule. Because the rule says the format should end up like a^2 - 2ab + b^2, and our equation began with x^2, I knew that a and x had to be the same. So I plugged x into a in the rule (so it looks like x^2 - 2bx + b^2). Then comparing the equations, you can see that they are different because of the 14 and the 49, right? You need to know what b is, so you set either 2bx=14 OR b^2=49, and solve. b should equal 7.

OpenStudy (anonymous):

I hope that makes sense. After you factor, you should get (x-7)^2. Now it's time to bring the 3 back in. Your equation should be (x-7)^2 = 3. Using the Square Root method just means take the square root of each side. If you square root (x-7)^2, you get x-7 and square-rooting 3 gives you \[\pm \sqrt{3}\]Now you solve for x. Add 7 to each side. You should then get 2 answers for x: \[x=7-\sqrt{3}\ OR\ x=7+\sqrt{3}\] in decimal form that's x=5.26795 OR x=8.73205. One of the answers matches one of the answer choices above.

OpenStudy (anonymous):

Ok thanx I have one more that I was wondering you if you could help me with?

OpenStudy (anonymous):

sure. but this post is getting kind of long. Wanna make a new post?

OpenStudy (anonymous):

ok